OpenStudy (anonymous):

Rationalize the denominator. 6/√(5 + √11)

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

my algebra is failing me on what to do after i multiply the bottom by rt 5

OpenStudy (anonymous):

i mean this is for a pre-cal review I'm doing

OpenStudy (anonymous):

yah but this concept is from alg 2

OpenStudy (anonymous):

I tried to multiply the bottom by sqrt 5-sqrt 11 but it didn't work

OpenStudy (anonymous):

too many square roots

OpenStudy (anonymous):

actually it is not clear what you wrote

OpenStudy (anonymous):

the answer is sqrt 11 - sqrt 5 but i need to know how to get the answer

OpenStudy (anonymous):

\[\frac{6}{\sqrt5+\sqrt{11}}\]?

OpenStudy (anonymous):

yes like that

OpenStudy (anonymous):

tats not what u wrote before

OpenStudy (anonymous):

multiply by the conjugate, you get it right away

OpenStudy (anonymous):

i tried :( would the conjugate be √5 - √11

OpenStudy (anonymous):

\[\frac{6}{\sqrt{5}+\sqrt{11}}\times \frac{\sqrt5-\sqrt{11}}{\sqrt{5}-\sqrt{11}}\] denominator is \(5-11=-6\) so you get \[\frac{6(\sqrt5-\sqrt{11})}{-6}\]

OpenStudy (anonymous):

then cancel you be done

OpenStudy (anonymous):

AH i forgot about conjugate

OpenStudy (anonymous):

its oposite the sign. thank you

OpenStudy (anonymous):

am i canceling 6 and -6?

OpenStudy (anonymous):

\[\frac{\cancel6(\sqrt5-\sqrt{11})}{-\cancel6}=-(\sqrt5-\sqrt{11})=\sqrt11-\sqrt5\]\]

OpenStudy (anonymous):

yes you are cancelling

OpenStudy (anonymous):

OHH i got it now thank you so much!

OpenStudy (anonymous):

yw